Make New Entry,
Make Followup Entry
User name J. Wexler
Log entry time 08:00:47 on August29,2009
Entry number 286766
keyword=Shift Summary - Owl 8/29
Shift summary - Owl 8/29:
12AM - Backing out of Moller tune.
12:50- In response to the target being cooled again, we start re-doing the carbon hole. Runs 24589-93, 4404-9.
1:07 - Spot++ has errors preventing us from finding position of the hole. Diagnosing.
1:40 - Moving on from spot++. Decide to go with 2 x 2.5 raster (4.4 x 4 MCC)
2:05 - Checked position of beam on BeO target before moving on to LH2 target.
2:25 - Ramping current on LH2 target.
2:50 - Reached 60 uA, but having FFB issues.
3:15 - DAQ issues, specifically panguin. Same target/current. Parity DAQ runs 11721-4.
4:55 - Shift checklist.
5:05 - Changed left arm HAPPEX trigger voltage to -1350. Also corrected pedestals. Run 11725-6.
5:15 - Zeroed beam to check pedestals, seem fine.
5:25 - Tell MCC we'd like to find PITA slopes.
5:35 - Begin taking data for PITA slopes. Run 11727.
5:50 - Begin analysis on PITA data. Take run 11728 while doing so, still on LH2 target, 60 uA, 2 x 2.5 raster.
6:20 - PITA scan data requires some editing, but the results look fine.
6:40 - Turn on IA feedback, put in PITA slopes. Runs 11729-30.
7:20 - Notice feedback on the IA.]
7:30 - Continuing to take data on the LH2. Run 11731.